At December 31, 2017, Crane Company had 1120000 shares of common stock outstanding. In addition, Crane had 455000 shares of preferred stock which were convertible into 759000 shares of common stock. During 2018, Crane paid $1250000 cash dividends on the common stock and $796000 cash dividends on the preferred stock. Net income for 2018 was $6800000 and the income tax rate was 40%. The diluted earnings per share for 2018 is (rounded to the nearest penny)

$2.55.
$3.62.
$6.06.
$5.40.
Answer #1

Basic earning per share = (6800000-796000)/1120000 = 5.36

Diluted earning per share = 6800000/(1120000+759000) = 3.62

So answer is b) $3.62